solve the undetermined coefficient y''+3y'+2y=e-x +e-2x -x

1
Expert's answer
2021-05-31T19:35:35-0400

Given the equation:


"y''+3y'+2y=e^{-x} +e^{-2x} -x"

We can rewrite the equation using the notation as:


"\frac{d^{2} y(x)}{d x^{2}}+3 \frac{d y(x)}{d x}+2 y(x)=e^{-x}+e^{-2 x}-x \\"

The general solution will be the sum of the complementary solution and particular solution. Find the complementary solution by solving:


"\frac{d^{2} y(x)}{d x^{2}}+3 \frac{d y(x)}{d x}+2 y(x)=0 :"

"\text{Assume a solution will be proportional to } e^{\lambda x} \text{ for some constant } \lambda.\\
\text{Substitute } y(x)=e^{\lambda x} \text{ into the differential equation:}\\"



"\frac{d^{2}}{d x^{2}}\left(e^{\lambda x}\right)+3 \frac{d}{d x}\left(e^{\lambda x}\right)+2 e^{\lambda x}=0\\
\text{Substitute } \frac{d^{2}}{d x^{2}}\left(e^{\lambda x}\right)=\lambda^{2} e^{\lambda x} \text{ and } \frac{d}{d x}\left(e^{\lambda x}\right)=\lambda e^{\lambda x}:\\
\lambda^{2} e^{x x}+3 \lambda e^{\lambda x}+2 e^{\lambda x}=0\\
\text{Factor out } e^{\lambda x} :
\left(\lambda^{2}+3 \lambda+2\right) e^{\lambda x}=0"

"\text{Since } e^{\lambda x} \neq 0 \text{for any finite }\lambda, \text{the zeros must come from the polynomial:}\\
\lambda^{2}+3 \lambda+2=0\\
Factor:\\
(\lambda+1)(\lambda+2)=0\\
Solve for \lambda :\\
\lambda=-2 or \lambda=-1\\
The root \lambda=-2 gives y_{1}(x)=c_{1} e^{-2 x} as a solution, where c_{1} is an arbitrary constant.\\

The root \lambda=-1 gives y_{2}(x)=c_{2} e^{-x} as a solution, where c_{2} is an arbitrary constant.\\
The general solution is the sum of the above solutions:\\
y(x)=y_{1}(x)+y_{2}(x)=c_{1} e^{-2 x}+c_{2} e^{-x}\\
Determine the particular solution to\\ \frac{d^{2} y(x)}{d x^{2}}+3 \frac{d y(x)}{d x}+2 y(x)=e^{-x}+e^{-2 x}-x \\\text{by the method of undetermined coefficients: }\\
The particular solution will be the sum of the particular solutions to\\
 \frac{d^{2} y(x)}{d x^{2}}+3 \frac{d y(x)}{d x}+2 y(x)=-x, \frac{d^{2} y(x)}{d x^{2}}+3 \frac{d y(x)}{d x}+2 y(x)=e^{-2 x} and\\
\frac{d^{2} y(x)}{d x^{2}}+3 \frac{d y(x)}{d x}+2 y(x)=e^{-x}"


"\text{ The particular solution to } \frac{d^{2} y(x)}{d x^{2}}+3 \frac{d y(x)}{d x}+2 y(x)=-x \text{ is of the form:}\\
y_{p_{1}}(x)=a_{1}+a_{2} x\\
\text{ The particular solution to }\frac{d^{2} y(x)}{d x^{2}}+3 \frac{d y(x)}{d x}+2 y(x)=e^{-2 x} is of the form:\\
y_{p_{2}}(x)=x\left(a_{3} e^{-2 x}\right), where a_{3} e^{-2 x} \text{ was multiplied by} x \text{ to account for }e^{-2 x} in the complementary solution.\\
The particular solution to \frac{d^{2} y(x)}{d x^{2}}+3 \frac{d y(x)}{d x}+2 y(x)=e^{-x} is of the form:\\
y_{p_{3}}(x)=x\left(a_{4} e^{-x}\right), where a_{4} e^{-x} was multiplied by x to account for e^{-x} in the complementary solution.\\
Sum y_{p_{1}}(x), y_{p_{2}}(x), and y_{p_{3}}(x) to obtain y_{p}(x) :\\
y_{p}(x)=y_{p_{1}}(x)+y_{p_{2}}(x)+y_{p_{3}}(x)=a_{1}+a_{2} x+a_{3} e^{-2 x} x+a_{4} e^{-x} x\\
Solve for the unknown constants a_{1}, a_{2}, a_{3}, and a_{4} :\\
Compute \frac{d y_{p}(x)}{d x}:\\
\frac{d y_{p}(x)}{d x}=\frac{d}{d x}\left(a_{1}+a_{2} x+a_{3} e^{-2 x} x+a_{4} e^{-x} x\right)\\
=a_{2}+a_{3} e^{-2 x}-2 a_{3} e^{-2 x} x+a_{4} e^{-x}-a_{4} e^{-x} x"


"Compute \frac{d^{2} y_{p}(x)}{d x^{2}}:\\
\frac{d^{2} y_{p}(x)}{d x^{2}}=\frac{d^{2}}{d x^{2}}\left(a_{1}+a_{2} x+a_{3} e^{-2 x} x+a_{4} e^{-x} x\right)\\
=a_{3}\left(-4 e^{-2 x}+4 e^{-2 x} x\right)+a_{4}\left(-2 e^{-x}+e^{-x} x\right)\\
Substitute the particular solution y_{p}(x) into the differential equation:\\
\frac{d^{2} y_{p}(x)}{d x^{2}}+3 \frac{d y_{p}(x)}{d x}+2 y_{p}(x)=e^{-x}+e^{-2 x}-x\\
a_{3}\left(-4 e^{-2 x}+4 e^{-2 x} x\right)+a_{4}\left(-2 e^{-x}+e^{-x} x\right)+
3\left(a_{2}+a_{3} e^{-2 x}-2 a_{3} e^{-2 x} x+a_{4} e^{-x}-a_{4} e^{-x} x\right)+
2\left(a_{1}+a_{2} x+a_{3} e^{-2 x} x+a_{4} e^{-x} x\right)=e^{-x}+e^{-2 x}-x\\
Simplify:\\
2 a_{1}+3 a_{2}-a_{3} e^{-2 x}+a_{4} e^{-x}+2 a_{2} x=e^{-2 x}+e^{-x}-x\\
Equate the coefficients of 1 on both sides of the equation:\\
2 a_{1}+3 a_{2}=0\\
Equate the coefficients of e^{-2 x} on both sides of the equation:\\
-a_{3}=1"


"Equate the coefficients of e^{-x} on both sides of the equation:\\
a_{4}=1\\
Equate the coefficients of x on both sides of the equation:\\
2 a_{2}=-1\\
Solve the system:\\
a_{1}=\frac{3}{4}\\
a_{2}=-\frac{1}{2}\\
a_{3}=-1\\
a_{4}=1\\
Substitute a_{1}, a_{2}, a_{3}, \text{ and }a_{4} into y_{p}(x)=a_{1}+x a_{2}+e^{-2 x} x a_{3}+e^{-x} x a_{4} :\\
y_{p}(x)=-\frac{x}{2}+e^{-x} x-e^{-2 x} x+\frac{3}{4}\\
The general solution is:\\

y(x)=y_{\mathrm{c}}(x)+y_{p}(x)=-\frac{x}{2}+e^{-x} x-e^{-2 x} x+c_{1} e^{-2 x}+c_{2} e^{-x}+\frac{3}{4}"


which is the required answer